One number is 15 more than the other. Find the numbers if their sum is 32.
Found 2 solutions by Cromlix, harpazo:
Answer by Cromlix(4381)

One number = x
Other number = 15 + x
x + (15 + x) = 32
2x + 15 = 32
2x = 32 - 15
2x = 17
x = 8.5
One number = 8.5
Other number = 23.5
Hope this helps.
:-)

Answer by harpazo(655)

x = 15 + y
x + y = 32
Use the substitution method to find your numbers.